OpenStudy (anonymous):

Find u_n, v_n, and t_N in terms of n for the following system: {u_n+1 = -u_n + 2v_n + t_n v_n+1 = v_n - t_n t_n+1 = 2t_n For u_0, v_0, and t_0 given "_" = subscripts EDIT: Intro to Linear Algebra

OpenStudy (anonymous):

The system \[u_{n+1} = -u_{n} + 2v_{n} + t_{n}\] \[v_{n+1} = v_{n} - t_{n}\] \[t_{n+1} = 2t_{n}\]

OpenStudy (anonymous):

Anyone? I really need help...